Find the heat required to heat 5-kg water from 20 °C to 80 °C. Express the temperature of 80 °C in both Fahrenheit degree and Kelvin
Answer #1

specific heat of water, S = 4186 J/(kg degree C )

for the heat needed to increase the temperature

heat needed = m * S* delta T

heat needed = 5 * 4186 * (80 - 20)

heat needed = 1255800 J

the heat needed to increase the temperature from 20 to 80 degree C is 1255800 J

-----

for 80 degree C

in kelvin

T = 80+ 273 = 353 K

the temperature is 353 K

to change to degree F

(T - 32)/180 = (80 - 0)/100

T = 176 degree F

the temperature is 176 degree F